Es provided results from a poll of 1500 aduhts who were asked to identity their aoe, Among the 1500 G: A
Answer #1

here, α=1-0.94=0.06

so, α/2 = 0.03

Z value at    0.03   =   1.881   (excel formula =NORMSINV(   0.03   ) )

so,Zα/2 = 1.88